OpenStudy (anonymous):

What is the slope of the graph of 8x - 4y = 11?

OpenStudy (anonymous):

Option A: -2 Option B: -½ Option C: ½ Option D: 2

OpenStudy (anonymous):

2

OpenStudy (mysesshou):

simplest to put into y=mx+b format 8x - 4y = 11 -4y=11-8x y=-11/4 + 2x y = 2x - 11/4 Yes, slope = 2